For each of the following expressions, simplify by expanding any brackets and collecting like terms. Show all the steps in your working.
(a) 4x +7 − 3x − 2
(b) 5(3 − 2x)
(c) 7(2y +1) − 3(4 − y)
(d) (N − 2)(2N +3)
(e) (4x − 3)^2
(f) 8-12p/4 + 4p

Answer by user_dude2008(1862)

a) 4x +7 - 3x - 2 = x + 5


b) 5(3 - 2x) = 15 - 10x


c) 7(2y +1) - 3(4 - y) = 14y+7 - 12+3y = 17y-5


d) (N-2)(2N+3) = 2N^2 + 3N - 4N - 6 = 2N^2 - N - 6


e) (4x - 3)^2 = (4x - 3)(4x - 3) = 16x^2 - 12x - 12x + 9 = 16x^2 - 24x + 9


f) (8-12p)/(4+4p) = (4(2-3p))/(4(1+p)) = (2-3p)/(1+p)